# Task-Agnostic Continual Prompt Tuning with Gradient-Based Selection and Decoding

### v1

@@ 4.2 Prompt Pool Compression via Gradient-Based Selection
where $\mu_{g}$ and $\sigma_{g}$ are the mean and standard deviation of gradient norms, and $\alpha$ is a tunable hyperparameter. We then define the subsets of the prompts as follows:
$\mathcal{P}_{high}=\{\mathbf{p}_{j}\mid g_{j}>\tau\},\quad\mathcal{P}_{low}=\{\mathbf{p}_{j}\mid g_{j}<\tau\}.\quad$ (3)
To prevent the complete loss of information encoded in low-gradient prompts $\mathcal{P}_{low}$, we propose an aggregation mechanism that consolidates them into a single aggregated prompt embedding $\mathbf{p}_{agg}$. We compute this aggregated prompt using gradient-weighted averaging as
@@
This weighting scheme ensures that prompts with relatively higher gradients—even among low-gradient ones—contribute more proportionally to the aggregated prompt. Consequently, the updated prompt pool for task $T_{t}$ is $\mathcal{P}^{\prime}=\mathcal{P}_{high}\cup\{\mathbf{p}_{agg}\}.$ This strategy retains essential knowledge while substantially reducing memory and inference costs. In particular, we observe minimal performance degradation even with a compressed prompt pool, indicating that less relevant prompts offer limited benefit for future tasks.
@@ 5.1 Datasets and Baselines
Datasets. Following Razdaibiedina [27], we evaluate our approach under two few-shot continual-learning settings: the “short-sequence” experiments and the “long-sequence” experiments. In addition, we introduce a Negative Transfer benchmark composed of sequences specifically chosen to induce transfer degradation, to assess robustness when task similarity is low. We also construct three long-sequence benchmark (order A-C) to analyze performance across different task difficulty progressions.
Long-Sequence CL Setting. We propose three novel task orders: (1) a Order $A$: easy-to-hard order, where tasks are categorically grouped and simpler tasks such as binary sentiment classification are introduced first, followed by more complex reasoning tasks like MultiRC and COPA; (2) a Order $B$: hard-to-easy order that reverses this progression; and (3) a Order $C$: mixed order that randomly interleaves tasks from different categories and difficulties. These variations allow us to assess how the order of task presentation and semantic similarity affect forgetting, transfer, and model generalization in long-horizon continual learning scenarios.
Negative Transfer Benchmark. To evaluate the robustness of our approach against negative transfer [24], we introduce a Negative Transfer Benchmark (NT1, NT2, NT3) inspired by the SHLPT framework [47]. Specifically, we identify task pairs that exhibit negative transfer based on the analysis provided in the SHLPT. Using these insights, we create three task sequences that are likely to cause negative transfer, meaning earlier tasks make it harder to learn later ones. This helps us test how well our method handles such challenges.
Baseline Methods for Comparison. In our experiments, we evaluate the proposed method against a total of 11 baseline methods222we could not compare with Q-Tuning[13] as they adopt a different setup and their codebase is not provided.: (1) Finetune [46], (2) Continual Finetune [46], (3) Prompt Tuning [18, 25], (4) Data Replay [8], (5) EWC [17], (6) A-GEM [6], (7) LFPT5 [25], (8) Per-task Prompt [18, 25], (9) ProgPrompt [27], (10) Coda-prompt [35], and (11) SHLPT [47].
We evaluate our method using multiple continual learning benchmarks. Specifically, we present the average test accuracy, backward transfer (BWT), and forward transfer (FWT) scores across different benchmark settings: short-sequence, long-sequence, easy-hard mixed, and negative transfer benchmark. All results are averaged over three runs for robustness. Further experimental metrics and details are provided in Appendix B.
(A) Baseline
(B) Ours
(C) B - A
Figure 2: Heatmaps of backward transfer scores on previous tasks for Order 4. (A) shows results from the baseline (Progressive Prompts), (B) shows our method, and (C) presents the difference (B - A). Brighter values indicate better retention of earlier tasks.
@@ 5.2 Backward Transfer Analysis
Table 3 presents the Backward Transfer (BWT) scores for long sequence, measuring how much the model forgets previous tasks while learning new ones. Our approach consistently reduces forgetting across all model variant compared to the baselines. Interestingly, although larger models generally achieve higher average accuracy (as seen in Table 12 in the Appendix), they tend to exhibit greater forgetting. In contrast, smaller models achieve better (i.e., less negative) BWT scores. One plausible explanation is that smaller models, with fewer parameters, may learn less aggressively on the current task, thereby generalizing better and retaining prior knowledge more effectively. This highlights a nuanced trade-off between model capacity and forgetting behavior in continual learning.
We observe consistent trends under the negative transfer benchmarks (NT1–NT3). As shown in Table 14 (Appendix), our method significantly reduces forgetting across all model sizes. Notably, Smaller models again show superior BWT, highlighting their robustness to task dissimilarity and the overall effectiveness of our approach in mitigating forgetting.
@@
To visualize BWT dynamics, Figures 2 and 3 offer detailed breakdowns. Figure 2 displays heatmaps of BWT scores for Order 4. Compared to the baseline (A), our method (B) exhibits brighter regions in the lower triangle, indicating stronger retention of earlier tasks. The difference heatmap (C) highlights widespread positive gains (blue), especially for early tasks, underscoring the effectiveness of compressing redundant prompts while preserving diverse, task-relevant knowledge. Figure 3 shows per-task BWT differences between our method and the baseline. Positive values indicate improved retention, while negative values signify degradation. Notably, our method yields significant BWT improvements on challenging tasks like boolq, sst2, and dbpedia_14, while maintaining parity or improvement on most others. This further confirms the effectiveness of our gradient-based compression in preserving essential task knowledge over long sequences.
One critical observation from our results is that Order B consistently outperforms Order A in both BWT scores and the number of forgotten tasks. Specifically, for the same set of tasks, Order B shows less negative BWT scores, indicating reduced forgetting. Additionally, the number of forgotten tasks is lower in Order B, further suggesting that harder tasks at the beginning of training help the model retain earlier knowledge more effectively. This observation highlights the benefit of starting with more difficult tasks, as it leads to better long-term retention and improved overall performance in continual learning. One also notes that Order C performs better than Order A but slightly less effectively than Order B, indicating that a mixed task sequence provides a balanced approach to task progression, offering a compromise between the benefits of easy-to-hard and hard-to-easy learning.
Figure 3: Per-task BWT comparison between our method (blue) and the baseline (red) for Order 4. Positive bars indicate improved retention of prior tasks. Our method shows significant BWT gains on several tasks (e.g., copa, wic, yahoo), demonstrating its effectiveness in mitigating forgetting across diverse task types.
Forgotten Task Count. We further analyze the number of forgotten tasks across various settings in Table 2. Our method significantly reduces the number of forgotten tasks compared to the Progressive Prompts baseline. For instance, in the case of T5-small, the number of forgotten tasks drops from as high as 93 in some orders to as low as 7 under our method. Similar improvements are observed for T5-base, where the count decreases from 90 to just 6 in the best case, and for T5-large, where it reduces from 87 to only 5. These improvements are consistent across all task orderings, including structured sequences such as Order A (easy-to-hard), Order B (hard-to-easy), and Order C (mixed). The results demonstrate that our gradient-based prompt compression effectively mitigates catastrophic forgetting, particularly in long-horizon continual learning setups, while preserving task-relevant knowledge in a compact form.
Table 2: Comparison of forgotten task counts between Progressive Prompts (Baseline) and our method for different models across various task orders (Order 4–6, A–C), now including mean forgotten tasks and relative improvement.
Model Setting Order 4 Order 5 Order 6 Order A Order B Order C Average Improvement (%)
T5-small ProgPrompt 85 85 88 90 84 93 87.5 76.8%
Ours 14 7 28 36 16 21 20.3
FT5-small ProgPrompt 93 95 95 94 97 93 94.5 76.7%
Ours 17 14 28 26 20 27 22.0
T5-base ProgPrompt 88 82 76 90 83 88 84.5 80.9%
Ours 15 6 19 29 13 15 16.2
FT5-base ProgPrompt 94 93 93 93 93 93 93.2 83.0%
Ours 10 10 15 25 14 21 15.8
T5-large ProgPrompt 77 72 80 87 71 85 78.7 82.4%
Ours 11 5 18 26 10 13 13.8
FT5-large ProgPrompt 92 93 93 93 91 93 92.5 79.8%
Ours 15 10 24 24 14 25 18.7
Table 3: Backward Transfer (BWT) comparison between our method and ProgPrompt across multiple task orderings (Order 4–6, A–C) for T5 and Flan-T5 models. Less negative scores indicate reduced forgetting. Our method consistently improves BWT across all model variants.
Model Setting Order 4 Order 5 Order 6 Order A Order B Order C Avg Imp. (%)
Ours -0.1402 -0.1894 -0.2504 -0.3141 -0.2734 -0.3006 -0.2447
Ours -0.0518 -0.1449 -0.2106 -0.2129 -0.1517 -0.2396 -0.1686
Ours -0.3060 -0.2404 -0.3549 -0.2893 -0.3439 -0.3259 -0.3101
Ours -0.2616 -0.3243 -0.3319 -0.3245 -0.3240 -0.2743 -0.3067
Ours -0.3243 -0.3336 -0.3979 -0.3912 -0.2956 -0.3512 -0.3490
Ours -0.3115 -0.3656 -0.3544 -0.3514 -0.3282 -0.3540 -0.3442
5.3 Forward Transfer Analysis
We also report FWT scores in Appendix B, showing that our improvements in BWT do not come at the cost of FWT. In fact, our method achieves comparable or even superior FWT in several cases, indicating that the prompt selection strategy not only mitigates forgetting but also avoids incorporating suboptimal prompts that could harm current task performance. These results highlight that enhancing BWT can be achieved without compromising forward transfer when prompts are selected judiciously.
5.4 Average Accuracy
Table 4: Average test-set accuracy on long-sequence experiments with the T5-large model.
Method T5-large
Order 4 Order 5 Order 6 Avg
Finetune 7.4 7.4 7.5 7.4
Prompt Tuning 8.2 8.7 7.9 10.9
Data Replay 55.2 54.8 53.3 54.4
LFPT5 70.4 68.2 69.1 69.2
Per-task Prompt 77.0 77.0 77.0 77.0
CODA-Prompt 76.2 77.0 76.4 76.5
ProgPrompt 75.7 78.6 74.3 76.2
SHLPT 77.6 76.9 77.8 77.4
Ours* 79.1 80.7 81.1 80.3
On the short-sequence setup (Table 20 in the Appendix), it attains 74.60% accuracy with T5-large, slightly outperforming ProgPrompt (74.5%) and surpassing baselines like LFPT5 and Continual Finetune. Since the short-sequence setup involves only four tasks, we retain all prompt candidates without compression. However, CODA-Prompt and SHLPT outperform our method slightly. These marginal differences are less significant given the simpler nature of short-sequence tasks. In contrast, our method excels in the long-sequence experiments (Table 4), achieving an average accuracy of 80.3%, surpassing all other baselines, including CODA-Prompt and SHLPT, demonstrating its superior performance in handling more complex, long-term task sequences. These results validate the effectiveness of our gradient-driven pruning approach in reducing redundancy while preserving essential task knowledge over long sequences. Detailed results across T5 and Flan-T5 model variants (Table 12 in the Appendix) show consistent gains, particularly for larger models (e.g., +3.2% for T5-large, +5.2% for FT5-base), highlighting the scalability and robustness of our method. Under the Negative Transfer benchmark (Orders NT1–NT3; Table 14 in the Appendix), our approach again outperforms all baselines, with improvements such as +3.8% on FT5-base and +3.6% on FT5-large, demonstrating strong generalization even in low-task-similarity scenarios.
@@ 5.5 Ablation Study
Table 5: Ablation study on the GRID framework over Orders 4–6, evaluating the contribution of gradient-based prompt selection (G) and task-aware decoding (D).
Model Variants 4 5 6 Avg
T5-large (0) G.R.I.D. -0.3243 -0.3336 -0.3979 –0.3519
(1) (w/o G) -0.3254 -0.3321 -0.3895 –0.3490
(1)+(2) w/o (D) -0.7032 -0.7589 -0.5967 –0.6863
(3) w/o G.R.I.D -0.7275 -0.7625 -0.6137 –0.7012
FT5-large (0) G.R.I.D. -0.3115 -0.3656 -0.3705 –0.3492
(1) (w/o G) -0.3145 -0.3510 -0.3723 –0.3459
(1)+(2) w/o (D) -0.7321 -0.7428 -0.6056 –0.6935
(3) w/o G.R.I.D -0.7679 -0.7444 -0.6195 –0.7106
To assess the contribution of each component in the GRID framework, we perform an ablation study on task Orders 4, 5 and 6. We evaluate the BWT scores under four configurations: (0) the complete G.R.I.D. model, (1) removing the gradient-based prompt selection (G), (2) removing task-aware decoding (D), and (3) removing both components (i.e., no GRID). As shown in Table 5, the full GRID model (row 0) achieves the best average BWT scores for both T5-large (-0.3519) and FlanT5-large (-0.3492), indicating strong resistance to forgetting. When only the gradient-based selection is removed (row 1), the performance remains nearly unchanged. The comparable scores confirm that prompt compression via gradient selection does not degrade retention. This is expected, as the goal of the gradient-based strategy is not to improve BWT directly, but to reduce prompt memory usage while maintaining similar performance.
Table 6: Prompt memory usage.
Method KB Reduction
ProgPrompt 600
Ours 200 66.7%
This memory advantage is quantitatively shown in Table 6, where our method requires only 200 KB of storage compared to 600 KB for ProgPrompt, achieving a 66.7% reduction. In contrast, removing the task-aware decoding component (row 2) leads to a noticeable drop in BWT, and removing both components (row 3) results in the worst performance for both models (-0.7012 for T5-large, -0.7106 for FlanT5-large). This highlights the critical role of decoding in backward knowledge retention, and the synergy between decoding and prompt selection in the full GRID setup. Overall, this ablation confirms that while the decoding component is central to improving BWT, the gradient-based prompt selection plays a complementary role by enabling scalability without compromising performance.
@@
Impact of Prompt Selection Strategies. Table 7 compares prompt selection strategies (Random, FIFO, and Gradient-based) on T5-large and FlanT5-large for Orders 4–6. Our gradient-based method consistently achieves the lowest forgotten task counts and competitive or best BWT scores across all settings. While FIFO occasionally matches our BWT, it retains more redundant prompts, leading to higher forgetting. Random selection performs worst overall, with unstable BWT and the highest forgetting. These results highlight the effectiveness of our approach in reducing memory while preserving prior knowledge.
@@
Table 7: Comparison of prompt selection strategies (Random, FIFO, Gradient-based) for T5-large and FlanT5-large across Orders 4–6. Metrics: average accuracy (Acc), BWT, and number of forgotten tasks (FT).
Model Strategy Order 4 Order 5 Order 6

### v3

@@ 3.2 Prompt Pool Compression via Gradient-Guided Selection
where $\mu_{g},\sigma_{g}$ are the mean and standard deviation of $\{g_{j}\}$, and $\alpha$ is a tunable hyperparameter:
$\mathcal{P}_{high}=\{\mathbf{p}_{j}:g_{j}>\tau\},\quad\mathcal{P}_{low}=\{\mathbf{p}_{j}:g_{j}<\tau\}.$ (3)
Although  equation 1 only measures prompt–task interaction, we observed that low-gradient prompts are often highly redundant (average cosine similarity $\geq 0.87$, Euclidean radius $R<0.45$). When similarity is high, discarding them has little effect; but when $\mathcal{P}_{low}$ is diverse, removal risks losing transferable information. To mitigate this, we aggregate $\mathcal{P}_{low}$ into a single embedding $\mathbf{p}_{agg}$ using gradient-weighted averaging:
@@
This ensures that relatively more informative low-gradient prompts contribute proportionally. The updated pool for task $T_{t}$ becomes $\mathcal{P}^{\prime}=\mathcal{P}_{high}\cup\{\mathbf{p}_{agg}\}.$
This mechanism preserves critical knowledge while substantially reducing memory and inference costs. Notably, experiments show minimal degradation under compression, confirming that low-gradient prompts contribute little to future tasks.
Why reduce storage? For every new task, all existing prompts are re-evaluated: only those with high gradient contributions are kept, while low-gradient or redundant ones are merged into a single aggregated prompt. This continual compression ensures that the pool size shrinks whenever redundancy is found. For example, after completing 15 tasks, ProgPrompt accumulates 15 prompts, whereas GRID retains 5, reducing storage by about 66.7% (Table 6). A smaller pool also decreases the number of prompts involved in forward/backward passes, which helps lower training time and can partly balance the additional gradient cost.
@@ 4.1 Datasets and Baselines
[…]
Baseline Methods for Comparison. We evaluate the proposed method against a total of 7 baseline methods [^1] We were unable to compare with Q-Tuning (Guo et al., 2024), since it follows a different setup and its codebase is not publicly available for reproduction in our setting due to its company privacy restrictions.: (1) Finetune (Wang et al., 2020), (2) Prompt Tuning (Lester et al., 2021; Qin & Joty, 2021), (3) Data Replay (de Masson D’Autume et al., 2019), (4) LFPT5 (Qin & Joty, 2021), (5) Per-task Prompt (Lester et al., 2021; Qin & Joty, 2021), (6) ProgPrompt (Razdaibiedina et al., 2023), and (7) SHLPT (Wu et al., 2024). We report average test accuracy, backward transfer (BWT), and forward transfer (FWT), defined by Lopez-Paz & Ranzato (2017) with results averaged over three runs; detailed metric definitions and additional results are provided in Appendix B.
4.2 Average Accuracy
Table 2: Average test-set accuracy on long-sequence order L1–L6 with T5-large. The DR column indicates whether the method uses data replay (✓) or not (✗).
Method L1 L2 L3 L4 L5 L6 DR Avg
Finetune 8.3 8.7 7.8 7.9 8.1 8.9 8.3
Prompt Tuning 8.8 9.5 8.1 9.3 9.4 9.4 9.1
Data Replay 56.2 54.3 53.5 54.8 54.2 55.3 54.7
LFPT5 70.8 69.2 69.4 68.2 69.4 68.5 69.3
Per-task Prompt 75.0 75.6 76.2 74.8 75.9 73.6 75.2
ProgPrompt 75.7 78.6 74.3 75.05 77.10 75.46 76.0
SHLPT 77.4 77.9 78.8 78.4 78.2 76.2 77.8
GRID* 79.1 80.7 81.1 79.0 79.8 75.5 79.2
Our method excels in the long-sequence experiments (L1-L3); Table 2), achieving an average accuracy of 79.2% and surpassing competitive baselines such as SHLPT, the most recent state-of-the-art method. These results validate the effectiveness of our gradient-driven pruning approach in reducing redundancy while preserving essential task knowledge. Detailed results across T5 and Flan-T5 variants (Table 13 in the Appendix) show consistent gains, particularly on larger models (e.g., +3.2% for T5-large, +5.2% for FT5-base). Under the Negative Transfer benchmarks (Order NT1–NT3; Table 16), we also observe clear improvements (e.g., +3.8% on FT5-base, +3.6% on FT5-large), highlighting its robustness even in low-task-similarity scenarios.
@@ 4.3 Backward Transfer Analysis
Table 3 and Table 11 (Appendix) report BWT scores across long-sequence and model variants. GRID consistently achieves substantially less negative BWT than ProgPrompt and SHLPT, cutting forgetting by nearly half on average. The gains hold across all T5 and Flan-T5 model sizes, with relative improvements exceeding 50%. Interestingly, while larger models achieve higher accuracy (Table 13), they suffer greater forgetting, whereas smaller models retain prior knowledge better, likely due to less aggressive adaptation to new tasks. To visualize BWT dynamics, Figure 2 displays heatmaps of BWT scores for Order L1. Compared to the progprompt (A) and SHLPT (B), our method (C) exhibits brighter regions in the lower triangle, indicating stronger retention of earlier tasks. The difference heatmap (D,E) highlights widespread positive gains (blue), especially for early tasks, underscoring the effectiveness of our approach. Figure 3 shows per-task BWT differences between our method and the baseline. Positive values indicate improved retention, while negative values signify degradation. Notably, our method yields significant BWT improvements on challenging tasks like boolq, sst2, and dbpedia_14, while maintaining parity or improvement on most others. Additional analysis of task order effects, including comparisons between Order L4, L5, and L6, is provided in Section  C.1, Tables 1314 and 17.
Table 3: BWT Score on T5-large across order (L1-L6) against the stronger baselines ProgPrompt and SHLPT. Less negative values indicate reduced forgetting.
Order ProgPrompt SHLPT GRID
L1 -0.7275 -0.6123 -0.3243
L2 -0.7625 -0.6870 -0.3336
L3 -0.6137 -0.5174 -0.3979
L4 -0.6257 -0.5042 -0.3912
L5 -0.6351 -0.4187 -0.2956
L6 -0.6416 -0.5840 -0.3512
Avg -0.6677 -0.5539 -0.3490
4.4 Forgotten Task Count
Table 4: Comparison of forgotten task counts across various task order L1-L6.
Order ProgPrompt SHLPT GRID
L1 77 64 11
L2 72 58 5
L3 80 69 18
L4 87 73 26
L5 71 58 10
L6 85 62 13
Avg 78.7 64.0 13.8
We define the Forgotten Task Count (FTC) as the number of tasks whose accuracy falls below a threshold relative to their standalone performance. Formally, task $T_{i}$ is forgotten at step $t$ if $a_{i}^{(t)}<\tau\cdot\min_{j}a_{j}^{(0)}$, where $a_{i}^{(0)}$ is the accuracy of $T_{i}$ when trained alone and $\tau\in(0,1)$. Using 0 as a cutoff is misleading since accuracies may degrade to small but nonzero values; our formulation instead provides a principled absolute threshold. [^2] In our task list, Amazon had the lowest standalone accuracy ($\sim$0.50); setting $\tau=0.4$ yields an absolute threshold of 0.20.. GRID reduces forgetting dramatically compared to both ProgPrompt and SHLPT. On average, GRID forgets only 13.8 tasks, compared to 78.7 for ProgPrompt and 64.0 for SHLPT. These results highlight that GRID mitigates catastrophic forgetting in long-horizon continual learning, while preserving task-relevant knowledge in a compact form. For results across different model variants, please refer to Table 14 in the Appendix.
Figure 2: Heatmaps of backward transfer scores on previous tasks for Order L1. (A) Progressive Prompts, (B) SHLPT, (C) GRID, and differences (D) C–A, (E) C–B.
4.5 Scalability to Larger Models
We have also run experiments on a larger model, T5-3B, to validate scalability. The results (Table 5) show that GRID maintains its advantages at this larger scale. These findings suggest that GRID’s core mechanisms scale effectively to models with billions of parameters.
Table 5: Comparison on T5-3B model
Order Method Accuracy BWT Forgot.
L1 ProgPrompt 74.86 -0.7161 69
SHLPT 77.18 -0.5841 57
GRID 78.54 -0.4445 21
L2 ProgPrompt 75.08 -0.7207 69
SHLPT 79.29 -0.6343 51
GRID 81.94 -0.5255 17
L3 ProgPrompt 74.14 -0.5840 81
SHLPT 79.56 -0.5632 61
GRID 81.85 -0.4152 33
@@ 4.6 Ablation Study
To evaluate the contribution of each component in the GRID framework, we conduct an ablation study on task order L1, L2, and L3, measuring BWT under four variants: (0) the full G.R.I.D. model, (1) without gradient-based prompt selection (G), (2) without constrained decoding (D), (3) without representative input selection (R), and (4) without all components. As shown in Table 6, the full model yields the highest BWT for both T5-large (-0.3519) and FlanT5-large (-0.3492). Excluding gradient-based selection has minimal effect on BWT, consistent with its role in reducing memory rather than directly improving retention: GRID requires only 200 KB of storage versus 600 KB for ProgPrompt, a 66.7% reduction. In contrast, removing constrained decoding causes a pronounced BWT drop, and eliminating both components leads to the worst performance (-0.7012 for T5-large, -0.7106 for FlanT5-large). Overall, these results underscore constrained decoding as the primary driver of backward transfer, with gradient-based selection offering complementary scalability.
Table 6: Ablation on GRID over orders L1–L3. We report BWT (less negative is better), average across orders, prompt memory size in KB (slots), and GPU time per run on A100 (40GB) with batch size 8.
Model Variant L1 L2 L3 Avg Memory GPU (h:m)
T5-large (0) G.R.I.D. 0.3243-0.3243 0.3310-0.3310 0.3979-0.3979 0.3511-0.3511 200200 27:08
(1) w/o G 0.3254-0.3254 0.3321-0.3321 0.3895-0.3895 0.3490-0.3490 600600 25:35
(2) w/o G,D 0.7032-0.7032 0.7589-0.7589 0.5967-0.5967 0.6863-0.6863 600600 26:42
(3) w/o G,D,R 0.7155-0.7155 0.7612-0.7612 0.5993-0.5993 0.6954-0.6954 600600 24:58
(4) w/o all 0.7275-0.7275 0.7625-0.7625 0.6137-0.6137 0.7012-0.7012 600600 23:58
FT5-large (0) G.R.I.D. 0.3115-0.3115 0.3456-0.3456 0.3705-0.3705 0.3425-0.3425 200200 27:10
(1) w/o G 0.3145-0.3145 0.3510-0.3510 0.3723-0.3723 0.3459-0.3459 600600 25:38
(2) w/o G,D 0.7321-0.7321 0.7428-0.7428 0.6056-0.6056 0.6935-0.6935 600600 26:37
(3) w/o G,D,R 0.7355-0.7355 0.7499-0.7499 0.6176-0.6176 0.7014-0.7014 600600 23:46
(4) w/o all 0.7679-0.7679 0.7444-0.7444 0.6195-0.6195 0.7106-0.7106 600600 23:47
Table 7: Accuracy for different datasets with varying sample sizes (k) highlights how model performance scales with more training samples.
k DBPedia Amazon AG News
20 0.5203 0.0000 0.0000
200 0.9674 0.0000 0.8180
1000 0.9880 0.5136 0.8900
2000 0.9892 0.5534
Runtime and Hardware. In addition to the A100 experiments in Table 6, we evaluated GRID on H100 GPUs (80GB) with batch size $16$. On H100, long task sequences averaged $\sim 11$ hours per run with peak memory usage of $\sim 65$ GB, compared to $\sim 27$ hours and $\sim 21$ GB on A100 (40GB, batch size $8$). These results demonstrate that GRID scales efficiently across hardware generations.
Role of Representative Inputs. We empirically find that increasing the number of representative samples beyond 1k per class leads to accuracy saturation with only negligible improvements. This indicates that larger sample sizes may not yield substantial performance gains (Table 7). By contrast, clustering provides both data efficiency and better generalization, highlighting the importance of representative input selection in GRID.
@@
Impact of Prompt Selection Strategies. As shown in Table 8, our gradient-based method consistently yields the fewest forgotten tasks and achieves competitive or superior BWT across all settings. FIFO occasionally matches our BWT but retains more redundant prompts, resulting in higher forgetting. Random selection performs worst, with unstable BWT and the largest number of forgotten tasks.
@@
Table 8: Comparison of prompt selection strategies (Random, FIFO, Gradient-based) for T5-large and FlanT5-large across order L1–L3. Metrics: average accuracy (Acc), BWT, and number of forgotten tasks (FT).
Model Strategy L1 L2 L3
@@
5 Related Work
Continual Learning. Continual learning (CL) involves learning from a sequence of tasks without full access to previous task data, aiming to preserve prior knowledge and enable positive transfer. A key challenge is catastrophic forgetting (McCloskey & Cohen, 1989), where updates to model parameters on new data erode earlier knowledge. CL strategies are categorized into three approaches: memory-based methods store and replay past task data to mitigate forgetting (Shin et al., 2017; Bang et al., 2021); regularization-based methods penalize deviations from important parameters to retain knowledge without stored data (Kirkpatrick et al., 2017; Zenke et al., 2017). Du et al. (2024) adopts a gradient-masking strategy by updating only high-activation model parameters, achieving task-agnostic and rehearsal-free CL; and architecture-based methods expand the model by adding new components for each task (Rusu et al., 2016; Yoon et al., 2018). These methods face scalability issues in large pretrained models, motivating the development of parameter-efficient CL techniques using lightweight components like prompts and adapters (Xu et al., 2023; Rücklé et al., 2021).
Continual Prompt Tuning. Prompt tuning (Lester et al., 2021; Li & Liang, 2021; Gu et al., 2021; Wang et al., 2023b; Jia et al., 2022) adapts large language models (LLMs) by learning a small set of continuous vectors, or soft prompts, prepended to the input tokens. Unlike full finetuning, it updates only the prompts while freezing model parameters, achieving competitive or superior performance with lower computational and memory cost. Continual prompt tuning (CPT) (Zhu et al., 2022; Yin et al., 2022; Ermis et al., 2022; Wang et al., 2022a) extends this idea to the continual learning (CL) setting, where models adapt to evolving task sequences. A substantial body of work has focused on improving CPT’s ability to retain and transfer knowledge, using techniques such as prompt concatenation (Razdaibiedina et al., 2023), parameter sharing (Wang et al., 2022b), and weighted prompt summation (Jiang et al., 2023). However, existing approaches often rely on large memory buffers to mitigate forgetting (Zhu et al., 2022; Ermis et al., 2022), impractical in privacy- or resource-constrained scenarios. Moreover, task-specific prompts suffer from latent forgetting under task-agnostic inference (Guo et al., 2024), while methods that expand prompt pools over tasks (Razdaibiedina et al., 2023; Wang et al., 2022a; Smith et al., 2023) face scalability and efficiency bottlenecks as tasks accumulate.
Gradient-Based Data and Parameter Selection. Gradient-based strategies have been extensively studied in data selection and coreset construction (Aljundi et al., 2019b), where the goal is to identify a representative or influential subset of training data (Borovicka et al., 2012; Rolf et al., 2021). Methods such as Coreset Selection (Killamsetty et al., 2021; Mirzasoleiman et al., 2020; Hao et al., 2023) and Gradient Matching (Aljundi et al., 2019b) leverage the similarity or norm of training gradients to preserve performance while reducing dataset sizes. In the context of prompt learning, some recent works have begun to reduce the prompt memory. For example, Q-Tuning (Guo et al., 2024) uses a PCA-based eviction and L2P (Wang et al., 2022b) keeps updating the same set of prompts over time. In contrast, our work draws inspiration from the coreset literature and extends it to the prompt space by dynamically evaluating gradient norms of prompt embeddings, allowing us to merge less informative prompts and improve scalability under long task sequences.